OpenStudy (anonymous):

Buckstars sells two types of coffee: one that costs $5.00 per pound and another that costs $6.20 per pound. How many pounds of the $5.00 per pound coffee should be mixed in with the $6.20 per pound coffee in order to have 30 pounds of a blend worth $5.50?

OpenStudy (anonymous):

the list will cost 5.00(x) where x = pounds. Then we have 30 -x. So 6.20(30-x). equation 5.00(x) + 6.20(30-x)= 30(5.50). 5.00x + 186- 6.20x= 165. then -1.2x=-21. Then x= 17.5 pounds.

OpenStudy (anonymous):

thank you :) <3333

OpenStudy (anonymous):

no problem :)